Status Report

NASA Space Station On-Orbit Status 14 June 2005

By SpaceRef Editor
June 15, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 14 June 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

Before breakfast, FE/SO Phillips deployed acoustic dosimeters on Sergei and himself, to be worn for 24 hours (with a microphone on the shirt collar). A third dosimeter was deployed in the Lab for a 24-hr. static data take (last time done: 5/11).  [Tonight, after about 15 hours of measurements, dosimeter data will be recorded and the hardware power-cycled, for another data take tomorrow morning after 8.5-hr. sleep. At that point, the crew will deploy the dosimeters statically in the station for the duration of the day, record measurements tomorrow night and stow the instruments. Acoustic data must be taken twice per Increment, each time for the duration of the 16-hour crew workday.]

John also deployed two passive FMK (formaldehyde monitoring kit) sampling assemblies in the Lab (below CEVIS) and Service Module (SM, most forward handrail), to catch any atmospheric formaldehyde on a collector substrate for subsequent laboratory analysis. (Last time done: 5/17).

The crew conducted final preparations for tomorrow’s undocking of Progress 17, transferring remaining Russian and US trash & disposable gear to the cargo ship with IMS (Inventory Management System) support and ground tagup.


Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

After stowing the last items for disposal and arranging them to maintain proper mass distribution for flight control equilibrium, Krikalev dismantled and removed the Progress LKT local temperature sensor commutator (TA251MB) of the BITS2-12 onboard telemetry system and its PZU-1M (ROM, read-only memory) unit, now no longer required and to be reused in the future. He also swapped a good SD1 lamp in the Progress with a bad (burnt-out) SD1 from the Service Module (SM).

Next, after activating 17P the crew disassembled the air duct in the hatchway to the SM PkhO (Transfer Compartment), then removed the threaded quick-disconnect (QD) screw clamps of the SM docking & internal transfer system (SSVP), which had rigidized the mating surfaces.  [The interface was visually inspected and video-recorded with the U.S. DVCAM to make sure that there is no damage to the cords, snap hooks or rings on the latches and to the slots for the clamps in the SSVP’s internal flanges. Before PrK/SU hatch closure, the interface video was downlinked via Ku-band for ground inspection.]

At ~2:15pm EDT, hatches between Progress and SM were closed, followed by depressurization of the transfer tunnel ( vestibule ) for the obligatory one-hour leak check.  [SM thrusters were disabled and automatic handover to the Russian segment (RS) was inhibited prior to the QD clamps removal. They were re-enabled after completed interface leak check. After completion of recharge of the Progress batteries, the 17P power system (RB) will be isolated from the SM power buses, providing the cargo ship with autonomous power for the undocking and post-undock flight phase. 17P will separate from the ISS tomorrow at 4:13pm EDT and reenter over the Pacific Ocean at ~7:51pm the same day.]

On the MedOps defibrillator equipment, John initiated the regular recharge process on battery #1 to full capacity and subsequently performed it also on battery #2. The task concluded with a battery voltage check.  [Each NiCad battery was charged for ~3.5 hrs, and its open-circuit voltage was tested at the end with the Aeolus volt/ohm scopemeter, then removed and stowed again. Nominally, the defib has a battery installed at all times, but with this particular unit the PDIM (power data interface module) is not functioning properly and would overcharge the batteries if left inside. They have to be charged every 60 days along with the defib checkout.]

The Science Officer set up the video equipment with the Lab camcorder to play back the first of two honey ops tapes from his recent (5/25) FMVM (Fluid Merging Viscosity Measurement) experiment for unattended downlink during Ku-band coverage.

Afterwards, the FE performed the monthly status check and filter cleaning, as required, on the autonomous PCG-STES010 (Protein Crystal Growth-Single Locker Thermal Enclosure System) payload in the Lab.  [For the PCG-STES, today is the 938th day of continuous powered operation onboard the ISS, which is unprecedented for this experiment hardware, which has delicate crystals growing inside. Its previous run time onboard Mir did not exceed 200 days. Nominally, PCG-STES010 powered operations are expected to continue while plans are finalized for its return on LF-1 (STS-114) next month.]

The FE handled the daily routine maintenance of the SM’s environment control & life support system (SOZh), including serving the toilet systems (ASU), and also prepared the regular IMS delta file for automated export/import to the IMS databases.

Both crewmembers worked out in their regular 2.5-hr. physical exercise program on the TVIS treadmill, RED resistive machine and VELO bike with bungee cord load trainer.  [Sergei s daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill and one hour on VELO plus load trainer (today: Day 2 of a new set).]

Afterwards, the FE transferred the exercise data files to the MEC for downlink, as well as the daily wristband HRM (heart rate monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

John also had time scheduled for the weekly TVIS maintenance in the current SLD (subject loading device) contingency configuration, primarily checking the condition of the SPDs (subject positioning devices) and recording time & date values.

At his discretion for today, Sergei had two tasks on his “time permitting” work list. The first was another session with the “Uragan” (hurricane) earth-imaging program, focusing the Nikon D1X No. 3 digital camera with 800-mm lens on targets called out on an uplinked list.  [Today’s targets included detailed strip photography with overlapping frames of Rostov Region fields and the lake system of the Kumo-Manych Depression, the Astrakhan game reserve located in a delta at the Caspian Sea, imagery of the entire area where the delta is in contact with the sea, the infrastructure of an oil production complex on the east coast of Caspian from the end point of Mangyshlak to Kara-Bogoz-Gol, plus scenes at crew s preference.]

As second “job jar” suggestion, the CDR conducted the regular daily inspection of the BIO-5 Rasteniya-2 experiment, including filling its water canister as required.  [Rasteniya (“plants”) researches growth and development of plants (currently horse radish) under spaceflight conditions in the Lada-7 greenhouse.]

At ~4:45am EDT, the Science Officer downlinked a message of greetings for replay to the participants of a Bone Loss During Spaceflight Symposium in Cleveland, OH, on 6/23.  [The two-day symposium will be attended by 25-30 scientists, and John was invited to ad-lib about his bone-preserving daily exercises and recent FOOT (Foot/Ground Reaction Forces during Space Flight) experiment session.]

At 2:30pm, John Phillips addressed members of the House Science Subcommittee at today s Hearing on Space and Aeronautics in a live TV downlink and audio uplink.  [Also testifying at the event were ISS Astronauts Peggy Whitson and Mike Fincke.]

Overnight, the ISS cabin atmosphere was repressurized with air from Progress 17 to raise ppN2 (nitrogen partial pressure) in preparation for the upcoming pressure equalization of PMA-3 (Pressurized Mating Adapter 3) on 6/20, which will result in a slight lowering of ppN2 (by about 8 mmHg).

Last night an RPC (remote power controller, #12) powering a heater on the MT (Mobile Transporter) tripped open in an external RPCM (RPC Module). This has happened before (12/25/2002) and is probably a Hybrid FET (field-effect transistor) failure, rather than a true overcurrent trip. A redundant RPC is operating, so for now RPC-12 has been left open and the Close command inhibited.

Update on Consumables:  As of yesterday (6/13), total oxygen (O2) on orbit is ~227 kg (~136 days). Progress 18, arriving on Saturday (6/18), will deliver an additional 85 days of O2, 96 days of water and 160 days of food.  [18P cargo: 110 kg O2, 40 new SFOG candles, 442 L water, 192 total rations of food in 64 food containers (21 US, 43 Russian).]

Update on Elektron: After the initial difficulty of transferring electrolyte (KOH, potassium hydroxide) from Liquid Unit 5 (BZh-5) to BZh-7 reported on 6/10, Sergei made a second attempt on Sunday (6/12), per uplinked ground instructions, and this time completed the transfer of 100-120 milliliter successfully. More KOH will be delivered on Progress-353/18P.

Update on Micropurification System (BMP): After the failure of the primary local digital commutator (LKTs #2V36) in the SM, reported yesterday, the BMP remains off since its telemetry data are questionable. TsUP/Moscow plans to switch to the backup LKTs in the next few days, which will allow BMP restart.  [The BMP is similar to the US TCCS (Trace Containment Control System).]

Today’s optional CEO (crew earth observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Dara Battlefield, Turkey (overlapping nadir mapping frames of the Dara, Turkey area are needed for archaeological investigations of an ancient battlefield. Vegetation differences in particular may help identify the location of a Roman trench and Persian siege mound in the region), Lake Nasser, Toshka Lakes, Egypt (water levels in the Toshka Lakes are expected to show fluctuations due to evaporation during the summer months. Water demands in L Nasser may also have an effect by reducing the supply of water to the Toshka water bodies. Nadir mapping along the shorelines of the lakes will be useful for assessment of water level fluctuations), Amazonian Fans, Brazil (weather is predicted to be mostly clear for nadir mapping along the orbit track. Looking for changes in vegetation and convoluted drainage networks that mark the boundaries of megafans), and Santa Barbara Coast, California (a coastal outflow was expected to produce clear conditions along the Santa Barbara coastline. These conditions may allow for photography of the islands located to the SW of the coastline that are included in the LTER site [Long Term Environmental Research, see 5/14 Status report]).

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 11 crew visit:

Expedition 11 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 6:47am EDT [= epoch]):

  • Mean altitude — 350.8 km
  • Apogee height — 353.5 km
  • Perigee height — 348.1 km
  • Period — 91.55 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0004015
  • Solar Beta Angle — 4.6 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.73
  • Mean altitude loss in last 24 hours — 200
  • Revolutions since FGB/Zarya launch (Nov. 98) — 37525

Some Increment 11 Main Events (not final):

  • Progress M-52 (17P) undock — 6/15 (4:13pm EDT);
  • Progress M-52 (17P) atmospheric entry — 6/15 (7.51pm EDT);
  • Progress M-53 (18P) launch — 6/16 (7:09pm EDT, Baikonur: 6/17, 5:09am)
  • Progress M-53 (18P) dock — 6/18 (8:44pm EDT);
  • PMA-3 depress — 6/22 (4:50am EDT);
  • Reboost — 6/29 (4:21pm, delta-V 2.3 m/s);
  • LF-1/STS-114 launch — NET 7/13 (18-day window opens);
  • LF-1/STS-114 dock — NET 7/15 (adding 110,713 kg to ISS mass);
  • Soyuz TMA-6 (10S) relocate (from DC-1 to FGB) — ~8/17;
  • Progress M-54 (19P) launch — 8/24 (dock 8/26);
  • ULF1.1/STS-121 launch — NET 9/9 (launch window opens);
  • 12A/STS-115 launch — NET 2/16/06;
  • 12A.1/STS-116 launch — NET 4/23/06;
  • 13A/STS-117 launch — NET 7/13/06.

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.